# Some Constructions on Quantum Principal Bundles

### v4

@@ 1. Universal Differential Envelope $\ast$–calculus in Differential Geometry
of $\mathbb{C}$–valued differential forms of $G$. As before, it suffices to consider convergent sequences $\{g_{i}\}^{\infty}_{i=1}$subseteq $H$ in $(\Gamma^{\wedge},d,\ast)$ to obtain $(C^{\infty}_{\mathbb{C}}(G)\otimes\bigwedge\mathfrak{g}^{\#}_{\mathbb{C}},d,\ast),$ although this lies beyond our purely geometric–algebraic approach.
The map $\Delta$ is a graded differential $\ast$–morphism. Thus, we have
$\Delta(g_{0}\,\wedge dg_{1}\wedge\cdots\wedge dg_{n})=\Delta(g_{0})\cdot d\Delta(g_{1})\cdots d\Delta(g_{n}).$
Moreover,
$\epsilon(g_{0}\,\wedge dg_{1}\wedge\cdots\wedge dg_{n})=\epsilon(g_{0})d\epsilon(g_{1})\cdots d\epsilon(g_{n})=0,$
as expected. Since $S$ is a graded antimultiplicative map that commutes with $d$, it follows that
$S(g_{0}\,\wedge dg_{1}\wedge\cdots\wedge dg_{n})=(-1)^{n(n-1)\over 2}dS(g_{n})\wedge\cdots\wedge dS(g_{1})\cdot S(g_{0}).$
@@
    $\Delta_{P}(x)=x\otimes\mathbbm{1}$ if and only if $x$in $B$.
@@
Note that if $\Delta_{\Omega^{\bullet}(P)}$ exists, it is unique because our graded differential $\ast$–algebras are generated by its degree 00 elements. Furthermore, $\Delta_{\Omega^{\bullet}(P)}$ is a graded differential $\Gamma^{\wedge}$–corepresentation on $\Omega^{\bullet}(P)$ [5]. In this setting, the space of horizontal forms is defined as
@@
[…]
@@
In accordance with [6], the set $\{\pi(z)\}$ is a left $P$–basis of $\Omega^{1}(P)$.
@@
[…]
@@
to coincide with $\Delta_{P}$ on the degree 00 case, and for degree $1$ elements,
@@
for all $g$in $P$. These constructions provide a differential calculus on $\zeta$. Consequently,
@@ 3. The Group Isomorphism
Let $\zeta$ be a qpb. Henceforth, we are going to consider that the quantum base space $(B,\cdot,\mathbbm{1},\ast)$ is a $C^{\ast}$–algebra (or it can be completed to a $C^{\ast}$–algebra). Under this assumption and in accordance with [8], for every $\delta^{V}$in $\mathcal{T}$ there exists
@@
for some $d_{V}$in $\mathbb{N}$ such that
@@
with $x^{V}_{ki}:=T^{\mathrm{L}}_{k}(e_{i})$, where $\mathcal{T}$ is a complete set of mutually non–equivalent irreducible (necessarily finite–dimensional) $\mathcal{G}$–corepresentations with $\delta^{\mathbb{C}}_{\mathrm{triv}}$in $\mathcal{T}$ (the trivial corepresentation on $\mathbb{C}$) and $\{e_{i}\}^{n_{V}}_{i=1}$ is the orthonormal basis of $V$ shown in Theorem 1.1.
@@
In particular, since $\delta^{\mathbb{C}}_{\mathrm{triv}}$in $\mathcal{T}$ we have that $\mathbbm{1}$in $\{g^{V}_{ij}\}_{\delta^{V},i,j}$ and
@@
when $\theta$in $\mathfrak{qg}^{\#}$, where
@@
for $\vartheta$in $H$, $\upsilon$in $\mathfrak{qg}^{\#}$ or $\vartheta$in $\mathfrak{qg}^{\#}$, $\upsilon$in $H$, with $\mathrm{qtrs}(\upsilon)=[\upsilon]_{1}\otimes_{\Omega^{\bullet}(B)}[\upsilon]_{2}$, where $\partial\vartheta$ is the grade of $\vartheta$. A direct calculation shows that the corresponding extended map $\widetilde{\mathrm{qtrs}}$ is the right inverse of $\widetilde{\beta}$ for degree $1$ and since $\widetilde{\beta}$ is bijective, we get that $\widetilde{\mathrm{qtrs}}$ is actually its inverse. It is worth mentioning that even when apparently the definition of $\mathrm{qtrs}$ depends on the real qpc $\omega$ chosen, by the uniqueness of the inverse function, the last result tells us that $\mathrm{qtrs}$ is independent of this choice. Since $\widetilde{\beta}$ commutes with the corresponding differential maps, it follows that ([10])
@@
for $g_{1}$,…, $g_{k}$in $H$. Now by considering the bilateral ideal $\mathcal{Q}$ generated by the relation presented on equation (14), a direct calculation shows that $\mathrm{qtrs}(\mathcal{Q})=0$ and therefore, $\mathrm{qtrs}$ can be defined on the universal differential envelope $\ast$–calculus
@@
    $[\vartheta]_{1}\,[\vartheta]_{2}=\epsilon(\vartheta)\mathbbm{1}$.
@@
    is the canonical graded twist map, i.e., $\sigma(\vartheta\otimes w)=(-1)^{kl}\,w\otimes\vartheta$ if $w$in $\Omega^{k}(P)$ and $\vartheta$in $\Gamma^{\wedge l}$.
@@
    $\mu\,\mathrm{qtrs}(\vartheta)=(-1)^{lk}\mathrm{qtrs}(\vartheta)\,\mu$ for all $\mu$in $\Omega^{k}(B)$, $\vartheta$in $\Gamma^{\wedge l}.$
@@
where $m_{\Omega^{\bullet}}:\Omega^{\bullet}(P)\otimes_{\Omega^{\bullet}(B)}\Omega^{\bullet}(P)\longrightarrow\Omega^{\bullet}(P)$ is the product map. We are going to show that $\mathfrak{f}_{\mathfrak{F}}$ is a convolution invertible map. First, since $\mathrm{qtrs}(\mathbbm{1})=\mathbbm{1}\otimes_{B}\mathbbm{1}$ it follows $\mathfrak{f}_{\mathfrak{F}}(\mathbbm{1})=\mathbbm{1}.$ Secondly, according to Proposition 3.3 point $3$
@@
is such that $\widehat{m}_{\Omega^{\bullet}}(w_{1}\otimes\vartheta_{1}\otimes_{\Omega^{\bullet}(B)}w_{2}\otimes\vartheta_{2})=(-1)^{kl}w_{1}w_{2}\otimes\vartheta_{1}\vartheta_{2}$ if $w_{2}$in $\Omega^{k}(P)$ and $\vartheta_{1}$in $\Gamma^{\wedge l}$. On the other hand, by equation (45) and Proposition 3.3 point $2$
@@
but by considering again Proposition 3.3 point $3$
@@
So for all $\vartheta$in $\Gamma^{\wedge}$
@@
Then for all $\vartheta$in $\Gamma^{\wedge}$

### v5

@@ 1. Universal Differential Envelope $\ast$–calculus in Differential Geometry
of $\mathbb{C}$–valued differential forms of $G$. As before, it suffices to consider convergent sequences $\{g_{i}\}^{\infty}_{i=1}$ $\subseteq$ $H$ in $(\Gamma^{\wedge},d,\ast)$ to obtain $(\Omega^{\bullet}_{\mathbb{C}}(G),d,\ast).$
Now we are going to proceed characterizing the graded differential $\ast$–Hopf algebra structure
$\Gamma^{\wedge\,\infty}=(\Gamma^{\wedge},\cdot,\mathbbm{1},\Delta,\epsilon,S,d,\ast).$
Let $\theta$ $\in$ $\mathfrak{g}^{\#}_{\mathbb{C}}$. Then $\theta=\pi(g)=(dg)_{e}$ for some $g$ $\in$ $H$. Taking $h$ $\in$ $H$, by equations (11), (18) and the fact that $\Delta$ is multiplicative, we have
$\Delta(h\,\pi(g))$ $=$ $h^{(1)}\otimes h^{(2)}\,\pi(g)+(h^{(1)}\otimes h^{(2)})\,\mathrm{ad}(\pi(g))$
$=$ $h^{(1)}\otimes h^{(2)}\,\pi(g)+h^{(1)}\,\pi(g^{(2)})\otimes h^{(2)}\,S(g^{(1)})g^{(3)}\;\in\;H\otimes\Gamma\oplus\Gamma\otimes H.$
Let $(A,(B,v))$ $\in$ $G\times TG_{\mathbb{C}}$, $((D,u),C)$ $\in$ $TG_{\mathbb{C}}\times G$ with $TG_{\mathbb{C}}=G\times\mathfrak{g}_{\mathbb{C}}$, for some $v$, $u$ $\in$ $\mathfrak{g}_{\mathbb{C}}$. Thus
$\Delta(h\,\pi(g))((A,(B,v))\oplus((D,u),C)$ $=$ $(h^{(1)}\otimes h^{(2)}\,\pi(g))(A,(B,v))$
$+$ $(h^{(1)}\,\pi(g^{(2)})\otimes h^{(2)}\,S(g^{(1)})g^{(3)})((D,u),C)$
$=$ $h^{(1)}(A)\,h^{(2)}(B)\,\pi(g)(v)$
$+$ $h^{(1)}(D)\,\pi(g^{(2)})(u)\,h^{(2)}(C)\,g^{(1)}(C^{-1})\,g^{(3)}(C)$
$=$ $h(AB)\,\pi(g)(v)$
$+$ $h(DC)\,g^{(1)}(C^{-1})\,\pi(g^{(2)})(u)\,g^{(3)}(C)$
$=$ $h(AB)\,\pi(g)(v)$
$+$ $h(DC)\,g^{(1)}(C^{-1})\,(dg^{(2)})_{e}(u)\,g^{(3)}(C)$
$=$ $h(AB)\,\pi(g)(v)$
$+$ $h(DC)\,\left.{d\,g(C^{-1}\,\mathrm{exp}(tu)\,C)\over dt}\right|_{t=0}.$
According to reference [15], the group product
$m_{TG}:=dm_{G}:TG_{\mathbb{C}}\times TG_{\mathbb{C}}\longrightarrow TG_{\mathbb{C}}$
of $TG_{\mathbb{C}}=G\times\mathfrak{g}_{\mathbb{C}}$ is given by
$m_{TG}((R,x),(S,y))=(RS,y+\mathrm{ad}^{\mathrm{class}}_{S}(x)),$
where $\mathrm{ad}^{\mathrm{class}}_{S}$ is the differential at $e$ of the classical right adjoint action $\mathrm{Ad}^{\mathrm{class}}_{S}$ by $S$ $\in$ $G$ on $G$, which is given by
$\mathrm{Ad}^{\mathrm{class}}:G\times G\longrightarrow G,\qquad\mathrm{Ad}^{\mathrm{class}}(R,S):=\mathrm{Ad}^{\mathrm{class}}_{S}(R):=S^{-1}RS;$
so
(39) $\mathrm{ad}^{\mathrm{class}}:\mathfrak{g}_{\mathbb{C}}\times G\longrightarrow\mathfrak{g}_{\mathbb{C}},\qquad\mathrm{ad}^{\mathrm{class}}(v,S):=\mathrm{ad}^{\mathrm{class}}_{S}(v)=(d\,\mathrm{Ad}^{\mathrm{class}}_{S})_{e}(v).$
Extending $m_{TG}$ to
[…]
we get
$m_{TG}(A,(B,v))=(AB,v),\qquad m_{TH}((D,u),C)=(DC,\mathrm{ad}^{\mathrm{class}}_{C}(u)).$
Thus
$((h\,\pi(g))\circ m_{TG})((A,(B,v))\oplus((D,u),C))$ $=$ $(h\,\pi(g))(AB,v)$
$+$ $(h\,\pi(g))(DC,\mathrm{ad}^{\mathrm{class}}_{C}(u))$
$=$ $h(AB)\,\pi(g)(v)$
$+$ $h(DC)\pi(g)(\mathrm{ad}^{\mathrm{class}}_{C}(u))$
$=$ $h(AB)\,\pi(g)(v)$
$+$ $h(DC)\,\left.{d\,g(C^{-1}\,\mathrm{exp}(tu)\,C)\over dt}\right|_{t=0}.$
Since every element $\vartheta$ of $\Gamma$ is the sum of elements of the form $h\,\pi(g)$, we conclude that
$\Delta(\vartheta)=\vartheta\circ m_{TG}.$
By equation (29), the last result and the facts that $(\Gamma^{\wedge},d,\ast)$ is generated by its degree 0 elements and $\Delta$ is a graded differential $\ast$–algebra morphism, it immediately follows that, for all $\vartheta$ $\in$ $\Gamma^{\wedge}$
(40) $\Delta(\vartheta)=\vartheta\circ m_{TG}.$
In light of last equation, we can extend the coproduct $\Delta:\Gamma^{\wedge}\longrightarrow\Gamma^{\wedge}\otimes\Gamma^{\wedge}$ to
(41) $\Delta:\Omega^{\bullet}_{\mathbb{C}}(G)\longrightarrow\Omega^{\bullet}_{\mathbb{C}}(G)\,\widetilde{\otimes}\,\Omega^{\bullet}_{\mathbb{C}}(G),\qquad\vartheta\longmapsto\vartheta\circ m_{TG},$
where $\widetilde{\otimes}$ denotes the corresponding completion of the algebraic tensor product $\otimes$ of graded differential $\ast$–algebras.
On the other hand, let $\pi(g)$ $\in$ $\mathfrak{g}^{\#}_{\mathbb{C}}$ and $h$ $\in$ $H$. In light of the definition of the antipode $S$ in equations (20), (31), we have
$S(h\,\pi(g))=S(\pi(g))\,S(h)$ $=$ $-\pi(g^{(2)})\,S(g^{3})\,S(S(g^{(1)}))\,S(h)$
$=$ $-\pi(g^{(2)})\,S(g^{3})\,g^{(1)}\,S(h)$
$=$ $-S(g^{3})\,g^{(1)}\,S(h)\,\pi(g^{(2)})\;\in\;\Gamma$
Let $(A,v)$ $\in$ $TG_{\mathbb{C}}=G\times\mathfrak{g}_{\mathbb{C}}$. Then
$S(h\,\pi(g))(A,v)$ $=$ $-(S(g^{3})\,g^{(1)}\,S(h)\,\pi(g^{(2)}))(A,v)$
$=$ $g^{(3)}(A^{-1})\,g^{(1)}(A)\,h(A^{-1})\,\pi(g^{(2)})(v)$
$=$ $-h(A^{-1})\,g^{(1)}(A)\,\pi(g^{(2)})(v)\,g^{(3)}(A^{-1})$
$=$ $-h(A^{-1})\,g^{(1)}(A)\,(dg^{(2)})_{e}(v)\,g^{(3)}(A^{-1})$
$=$ $-h(A^{-1})\,\left.{d\,g(A\,\mathrm{exp}(tv)\,A^{-1})\over dt}\right|_{t=0}.$
According to [15], the map
$\iota_{TG}:=d\iota_{G}:TG_{\mathbb{C}}\longrightarrow TG_{\mathbb{C}}$
that sends an element of $TG_{\mathbb{C}}=G\times\mathfrak{g}_{\mathbb{C}}$ to its multiplicative inverse, is given by
$\iota_{TG}(A,v)=(A,v)^{-1}=(A^{-1},-\mathrm{ad}^{\mathrm{class}}_{A^{-1}}(v)).$
In this way, we have
$((h\,\pi(g))\circ\iota_{TG})(A,v)$ $=$ $(h\,\pi(g))(A^{-1},-\mathrm{ad}^{\mathrm{class}}_{A^{-1}}(v))$
$=$ $-h(A^{-1})\,\pi(g)(\mathrm{ad}^{\mathrm{class}}_{A^{-1}}(v))$
$=$ $-h(A^{-1})\,\left.{d\,g(A\,\mathrm{exp}(tv)\,A^{-1})\over dt}\right|_{t=0}.$
Since every element $\vartheta$ of $\Gamma$ is the sum of elements of the form $h\,\pi(g)$, we conclude that
$S(\vartheta)=\vartheta\circ\iota_{TG}.$
By equation (31), the last result and the facts that $(\Gamma^{\wedge},d,\ast)$ is generated by its degree 0 elements and $S$ is a graded antimultiplicative linear map that commutes with the differential, it immediately follows that, for all $\vartheta$ $\in$ $\Gamma^{\wedge}$
(42) $S(\vartheta)=\vartheta\circ\iota_{TG}.$
In light of last equation, we can extend the coantipode $S:\Gamma^{\wedge}\longrightarrow\Gamma^{\wedge}$ to
(43) $S:\Omega^{\bullet}_{\mathbb{C}}(G)\longrightarrow\Omega^{\bullet}_{\mathbb{C}}(G),\qquad\vartheta\longmapsto\vartheta\circ\iota_{TG}.$
Finally, according to [15], the multiplicative neutral element of $TG_{\mathbb{C}}=G\times\mathfrak{g}_{\mathbb{C}}$ is $e\cong(e,0)$; thus (see equation (19))
$\vartheta(e,0)=0=\epsilon(\vartheta)$
for all $\vartheta$ $\in$ $\Gamma$. By the last result and the facts that $(\Gamma^{\wedge},d,\ast)$ is generated by its degree 0 elements and $\epsilon$ is a graded differential $\ast$–algebra morphism, it immediately follows that
$\epsilon:\Gamma^{\wedge}\longrightarrow\mathbb{C}$
is the evaluation at $e$ $\in$ $G$. As before, we can extend the counit $\epsilon$ to
(44) $\epsilon:\Omega^{\bullet}_{\mathbb{C}}(G)\longrightarrow\Omega^{\bullet}_{\mathbb{C}}(G),\qquad\vartheta\longmapsto\vartheta(e).$
Hence, we conclude that the structure of graded differential $\ast$–Hopf algebra of $\Gamma^{\wedge}$ reflects the fact that for a given Lie group, its complexified tangent bundle $TG_{\mathbb{C}}=G\times\mathfrak{g}_{\mathbb{C}}$ is also a Lie group ([15]).
It is worth mentioning that the previous calculations also show that the (right) adjoint HH–coaction on $H$ (see equation (12)) is the pull–back of the classical (right) adjoint action of $G$ on $G$, and the (right) adjoint HH–coaction on $\mathfrak{qh}^{\#}=\mathfrak{g}^{\#}_{\mathbb{C}}$ (see equation (10)) is the pull–back of the classical (right) adjoint action of $G$ on $\mathfrak{g}_{\mathbb{C}}$. In other words
(45) $\mathrm{Ad}(g)=g\circ\mathrm{Ad}^{\mathrm{class}}$
for all $g$ $\in$ $H$, and
(46) $\mathrm{ad}(\theta)=\theta\circ\mathrm{ad}^{\mathrm{class}}$
for all $\theta$ $\in$ $\mathfrak{g}^{\#}_{\mathbb{C}}$.
Consider the classical adjoint Lie algebra representation of $\mathfrak{g}_{\mathbb{C}}$
(47) $\mathrm{c}:\mathfrak{g}_{\mathbb{C}}\times\mathfrak{g}_{\mathbb{C}}\longrightarrow\mathfrak{g}_{\mathbb{C}}$
which is given by
$\mathrm{c}(v,u)$ $=$ $\left.{d\,\mathrm{ad}^{\mathrm{class}}_{\mathrm{exp}(tu)}(v)\over dt}\right|_{t=0}$
$=$ $\left.{d\,(d\,\mathrm{Ad}^{\mathrm{class}}_{\mathrm{exp}(tu)})_{e}(v)\over dt}\right|_{t=0}$
$=$ $\left.{d\over dt}\left(\left.{d\over dx}(\mathrm{exp}(tu)^{-1}\,\mathrm{exp}(xv)\,\mathrm{exp}(tu))\right|_{x=0}\right)\right|_{t=0}$
$=$ $[v,u]_{\mathbb{C}},$
where $[-,-]_{\mathbb{C}}$ is the complexified Lie bracket.
Now, consider the linear map
(48) $\mathrm{c}^{T}:=(\mathrm{id}_{\mathfrak{g}^{\#}_{\mathbb{C}}}\otimes\pi)\circ\mathrm{ad}:\mathfrak{g}^{\#}_{\mathbb{C}}\longrightarrow\mathfrak{g}^{\#}_{\mathbb{C}}\otimes\mathfrak{g}^{\#}_{\mathbb{C}}$
given by
$\mathrm{c}^{T}(\theta)=\mathrm{c}^{T}(\pi(g))=\pi(g^{(2)})\otimes\pi(S(g^{(1)})g^{(3)})$
for all $\theta$ $\in$ $\mathfrak{g}^{\#}_{\mathbb{C}}$.
Then, for all $v$, $u$ $\in$ $\mathfrak{g}_{\mathbb{C}}$, we have
$(\mathrm{c}^{T}(\theta))(v,u)$ $=$ $(\pi(g^{(2)})\otimes\pi(S(g{(1)})g^{(3)}))(v,u)$
$=$ $\pi(g^{(2)})(v)\,\pi(S(g{(1)})g^{(3)})(u)$
$=$ $\left(\left.{dg^{(2)}(\mathrm{exp}(xv))\over dx}\right|_{x=0}\right)\left(\left.{d\over dt}(g^{(1)}(\mathrm{exp}(tu)^{-1})\,g^{(3)}(\mathrm{exp}(tu))\right|_{t=0}\right)$
$=$ $\left.{d\over dt}(g^{(1)}(\mathrm{exp}(tu)^{-1})\,\left.{dg^{(2)}(\mathrm{exp}(xv))\over dx}\right|_{x=0}\,g^{(3)}(\mathrm{exp}(tu))\right|_{t=0}$
$=$ $\left.{d\over dt}(g^{(1)}(\mathrm{exp}(tu)^{-1})\,(dg^{(2)})_{e}(v)\,g^{(3)}(\mathrm{exp}(tu))\right|_{t=0}$
$=$ $\left.{d\over dt}\left(\left.{d\over dx}g(\mathrm{exp}(tu)^{-1}\,\mathrm{exp}(xv)\,\mathrm{exp}(tu))\right|_{x=0}\right)\right|_{t=0}$
$=$ $(dg)_{e}\left(\left.{d\over dt}\left(\left.{d\over dx}(\mathrm{exp}(tu)^{-1}\,\mathrm{exp}(xv)\,\mathrm{exp}(tu))\right|_{x=0}\right)\right|_{t=0}\right)$
$=$ $(dg_{e})(c(v,u))$
$=$ $\pi(g)(c(v,u))=\theta(c(v,u)).$
We conclude that
(49) $\mathrm{c}^{T}(\theta)=\theta\circ\mathrm{c}.$
for all $\theta$ $\in$ $\mathfrak{g}^{\#}_{\mathbb{C}}$.
The example developed explicitly shows that the universal differential envelope $\ast$–calculus $(\Gamma^{\wedge},d,\ast)$ is a proper generalization of the algebra $(\Omega^{\bullet}_{\mathbb{C}}(G),d,\ast)$ of $\mathbb{C}$–valued differential forms of $G$ in non–commutative geometry. In this way, for a given quantum group $\mathcal{G}$ and a bicovariant $\ast$–FODC $(\Gamma,d)$ over a quantum group $\mathcal{G}$, the triplet
$(\Gamma^{\wedge},d,\ast)$
will be interpreted as the $\ast$–algebra of quantum differential forms of $\mathcal{G}$. In this sense, the quantum dual Lie algebra
$\mathfrak{qh}^{\#}={\mathrm{Ker}(\epsilon)\over\mathcal{R}}$
plays the role of the dualization $\mathfrak{g}^{\#}_{\mathbb{C}}$ of $\mathfrak{g}_{\mathbb{C}}$ and the HH–corepresentation $\mathrm{ad}$ plays the role of the dualization of the right adjoint action $\mathrm{ad}^{\mathrm{class}}$ of $G$ on $\mathfrak{g}_{\mathbb{C}}$ (see equation (46)). Furthermore, by equation (49), the linear map
(50) $\mathrm{c}^{T}:=(\mathrm{id}_{\mathfrak{qh}^{\#}}\otimes\pi)\circ\mathrm{ad}:\mathfrak{qh}^{\#}\longrightarrow\mathfrak{qh}^{\#}\otimes\mathfrak{qh}^{\#}$
will be interpreted as the quantum Lie bracket.
@@
    $\Delta_{P}(x)=x\otimes\mathbbm{1}$ if and only if $x$ $\in$ $B$.
@@
Note that if $\Delta_{\Omega^{\bullet}(P)}$ exists, it is unique because our graded differential $\ast$–algebras are generated by its degree 0 elements. Furthermore, $\Delta_{\Omega^{\bullet}(P)}$ is a graded differential $\Gamma^{\wedge}$–corepresentation on $\Omega^{\bullet}(P)$ [5]. In this setting, the space of horizontal forms is defined as
@@
be its dense $\ast$–Hopf algebra. Consider $B:=\{\lambda\,\mathbbm{1}\mid\lambda\in\mathbb{C}\}$ and define the triplet
@@
In accordance with [6], the set $\{\pi(z)\}$ is a left PP–basis of $\Omega^{1}(P)$.
@@
In this case, $\Gamma=H\otimes\{0\}=\{0\}$, so $\Gamma^{\wedge}=\Gamma^{\wedge\,0}=H$, and $d=0$.
@@
to coincide with $\Delta_{P}$ on the degree 0 case, and for degree $1$ elements,
@@
for all $g$ $\in$ $P$. These constructions provide a differential calculus on $\zeta$. Consequently,
@@ 3. The Group Isomorphism
Let $\zeta$ be a qpb. Henceforth, we are going to consider that the quantum base space $(B,\cdot,\mathbbm{1},\ast)$ is a $C^{\ast}$–algebra (or it can be completed to a $C^{\ast}$–algebra). Under this assumption and in accordance with [8], for every $\delta^{V}$ $\in$ $\mathcal{T}$ there exists
@@
for some $d_{V}$ $\in$ $\mathbb{N}$ such that
@@
with $x^{V}_{ki}:=T^{\mathrm{L}}_{k}(e_{i})$, where $\mathcal{T}$ is a complete set of mutually non–equivalent irreducible (necessarily finite–dimensional) $\mathcal{G}$–corepresentations with $\delta^{\mathbb{C}}_{\mathrm{triv}}$ $\in$ $\mathcal{T}$ (the trivial corepresentation on $\mathbb{C}$) and $\{e_{i}\}^{n_{V}}_{i=1}$ is the orthonormal basis of $V$ shown in Theorem 1.1.
@@
In particular, since $\delta^{\mathbb{C}}_{\mathrm{triv}}$ $\in$ $\mathcal{T}$ we have that $\mathbbm{1}$ $\in$ $\{g^{V}_{ij}\}_{\delta^{V},i,j}$ and
@@
when $\theta$ $\in$ $\mathfrak{qg}^{\#}$, where
@@
for $\vartheta$ $\in$ $H$, $\upsilon$ $\in$ $\mathfrak{qg}^{\#}$ or $\vartheta$ $\in$ $\mathfrak{qg}^{\#}$, $\upsilon$ $\in$ $H$, with $\mathrm{qtrs}(\upsilon)=[\upsilon]_{1}\otimes_{\Omega^{\bullet}(B)}[\upsilon]_{2}$, where $\partial\vartheta$ is the grade of $\vartheta$. A direct calculation shows that the corresponding extended map $\widetilde{\mathrm{qtrs}}$ is the right inverse of $\widetilde{\beta}$ for degree $1$ and since $\widetilde{\beta}$ is bijective, we get that $\widetilde{\mathrm{qtrs}}$ is actually its inverse. It is worth mentioning that even when apparently the definition of $\mathrm{qtrs}$ depends on the real qpc $\omega$ chosen, by the uniqueness of the inverse function, the last result tells us that $\mathrm{qtrs}$ is independent of this choice. Since $\widetilde{\beta}$ commutes with the corresponding differential maps, it follows that ([10])
@@
for $g_{1}$,…, $g_{k}$ $\in$ $H$. Now by considering the bilateral ideal $\mathcal{Q}$ generated by the relation presented on equation (15), a direct calculation shows that $\mathrm{qtrs}(\mathcal{Q})=0$ and therefore, $\mathrm{qtrs}$ can be defined on the universal differential envelope $\ast$–calculus
@@
    $\vartheta]_{1}\,[\vartheta]_{2}=\epsilon(\vartheta)\mathbbm{1}$.
@@
    is the canonical graded twist map, i.e., $\sigma(\vartheta\otimes w)=(-1)^{kl}\,w\otimes\vartheta$ if $w$ $\in$ $\Omega^{k}(P)$ and $\vartheta$ $\in$ $\Gamma^{\wedge l}$.
@@
    $\mu\,\mathrm{qtrs}(\vartheta)=(-1)^{lk}\mathrm{qtrs}(\vartheta)\,\mu$ for all $\mu$ $\in$ $\Omega^{k}(B)$, $\vartheta$ $\in$ $\Gamma^{\wedge l}.$
@@
where $m_{\Omega^{\bullet}}:\Omega^{\bullet}(P)\otimes_{\Omega^{\bullet}(B)}\Omega^{\bullet}(P)\longrightarrow\Omega^{\bullet}(P)$ is the product map. We are going to show that $\mathfrak{f}_{\mathfrak{F}}$ is a convolution invertible map. First, since $\mathrm{qtrs}(\mathbbm{1})=\mathbbm{1}\otimes_{B}\mathbbm{1}$ it follows $\mathfrak{f}_{\mathfrak{F}}(\mathbbm{1})=\mathbbm{1}.$ Secondly, according to Proposition 3.3 point 33
@@
is such that $\widehat{m}_{\Omega^{\bullet}}(w_{1}\otimes\vartheta_{1}\otimes_{\Omega^{\bullet}(B)}w_{2}\otimes\vartheta_{2})=(-1)^{kl}w_{1}w_{2}\otimes\vartheta_{1}\vartheta_{2}$ if $w_{2}$ $\in$ $\Omega^{k}(P)$ and $\vartheta_{1}$ $\in$ $\Gamma^{\wedge l}$. On the other hand, by equation (70) and Proposition 3.3 point 22
@@
but by considering again Proposition 3.3 point 33
@@
So for all $\vartheta$ $\in$ $\Gamma^{\wedge}$
@@
Then for all $\vartheta$ $\in$ $\Gamma^{\wedge}$